# EVF-SAM: Early Vision-Language Fusion for Text-Prompted Segment Anything Model

### v4

@@ Abstract
Segment Anything Model (SAM) has attracted widespread attention for its superior interactive segmentation capabilities with visual prompts while lacking further exploration of text prompts. In this paper, we empirically investigate what text prompt encoders (e.g., CLIP or LLM) are good for adapting SAM for referring expression segmentation and introduce the Early Vision-language Fusion-based SAM (EVF-SAM). EVF-SAM is a simple yet effective referring segmentation method which exploits multimodal prompts (i.e., image and text) and comprises a pre-trained vision-language model to generate referring prompts and a SAM for segmentation. Surprisingly, we observe that: (1) multimodal prompts and (2) vision-language models with early fusion (e.g., BEIT-3) are beneficial for prompting SAM for accurate referring segmentation. Our experiments show that the proposed EVF-SAM based on BEIT-3 can obtain state-of-the-art performance on RefCOCO/+/g for referring expression segmentation and demonstrate the superiority of prompting SAM with early vision-language fusion. In addition, the proposed EVF-SAM with 1.32B parameters achieves remarkably higher performance while reducing nearly 82% of parameters compared to previous SAM methods based on large multimodal models. Code and models will be made publicly available.
@@ 1 Introduction
Segment Anything Model (SAM) (Kirillov et al., 2023) brings interactive segmentation paradigm to public view. Well-trained on the SA-1B dataset, SAM achieves stunning performance and quickly becomes popular as a vision foundation model for object localization and beyond. Various SAM variants (Xiong et al., 2023; Zhang et al., 2023; Zhao et al., 2023; Ke et al., 2024) have been explored, achieving better efficiency or higher precision. Despite SAM’s surprising abilities like point-prompted and box-prompted segmentation, it is a pity that the text-prompted segmentation ability remains conceptual. We retrospect such task to Referring Expression Segmentation (RES). RES focuses on the solution that one predicts the segmentation mask according to the text description given by users, which enjoys several explorations by some traditional end-to-end models (Hu et al., 2016; Liu et al., 2017; Shi et al., 2018; Chen et al., 2019; Ye et al., 2019; Hu et al., 2020; Ding et al., 2021; Li & Sigal, 2021b; Wang et al., 2022b; Yang et al., 2022; Liu et al., 2023c; Wu et al., 2023; Liu et al., 2023e; Yan et al., 2023), and is broadened by some Large Multimodal Models (LMM) (Lai et al., 2023; Yang et al., 2023; Ren et al., 2023; Pi et al., 2023; Xu et al., 2023; Zhang et al., 2024; Xia et al., 2023; Rasheed et al., 2023).  [^1] Equal contribution.  [^2] Corresponding author (xgwang@hust.edu.cn).
Figure 2: Comparisons of different Text-prompted SAM. (a) Given input texts, several works (Ren et al., 2024) leverage grounded detectors, e.g., Grounding DINO (Liu et al., 2023d), to generate box prompts for SAM. (b) A natural idea to support text prompts is to use an off-the-shelf text encoder to generate text embeddings for SAM (Kirillov et al., 2023; Li et al., 2023) while the performance of referring segmentation is inferior. (c) Several works (Lai et al., 2023; Yang et al., 2023; Rasheed et al., 2023) adopt Large Language Models (LLM) or Large Multimodal Models (LMM) to generate prompt embeddings for SAM in an autoregressive manner, which incurs a large computation burden. (d) Our proposed EVF-SAM exploits an effective Multimodal Encoder for text-prompted SAM with higher performance and fewer parameters compared to LLM-based methods.
The key challenge lies in empowering SAM with language understanding ability for segmentation according to text prompts, e.g., referring expression segmentation. Fig. 2 summarizes previous works which explore the text-prompted abilities of SAM: (a) SAM with grounded detector: A two-stage framework where a grounded detector generates a bounding box to prompt SAM, e.g., Grounded-SAM (Ren et al., 2024). However, those methods suffer from a sub-optimal architecture, where segmentation heavily relies on the accuracy of the detector, and it is difficult to optimize due to its non-end-to-end nature. (b) SAM with text encoder: A off-the-shelf text encoder, e.g., CLIP (Radford et al., 2021), is used to encode the text prompt, providing text embeddings for SAM. Whereas the semantic gap exists between the text embeddings and SAM which is pre-trained with geometric prompts, i.e., points or boxes, thus the segmentation performance is inferior. (c) SAM with LLM: A Large Language Model (LLM) (or Large Multimodal Model) is employed and fine-tuned to get the desired embeddings about object information. The embeddings will be used to predict segmentation masks based on image features. However, these LLM-based models are often computationally expensive, requiring massive memory and computation budgets, and the training is challenging. Additionally, complex conversation templates need to be manually designed to instruct the LLM for referring segmentation. Can we leverage a more efficient but effective method to empower SAM with text-prompted ability in an end-to-end manner?
To this end, we empirically investigate how to encode text prompts for SAM to address referring expression segmentation. Interestingly, we observe that (1) using multimodal prompts including both the text and image performs better than the text-only prompts and (2) the Multimodal Encoders with early vision-language fusion demonstrate significant superiority compared to text-only encoders or Large Language Models, as shown in Fig. 2 (d).
Motivated by the above observations, we extend SAM for language understanding and text-prompt capabilities by incorporating a Multimodal Encoder with Early Vision-Language Fusion (EVF) and present EVF-SAM in this paper. The proposed EVF-SAM aims to be a simple framework to prompt SAM with texts and illustrate how to prompt SAM to follow referring expressions effectively. EVF-SAM is built on the off-the-shelf foundation models and comprises a Multimodal Encoder, an early-fused vision-language model, e.g., BEIT-3 (Wang et al., 2022a), and a simple projector to generate prompt embeddings for SAM. EVF-SAM does not include elaborate designs or modules and is easy for scaling to larger models.
Training EVF-SAM is simple and conducted on referring segmentation datasets, e.g., RefCOCO (Yu et al., 2016), which is appropriate to adapt the original SAM for text prompts. Despite the simple architecture, our EVF-SAM achieves superior performance on referring expression segmentation tasks and outperforms previous attempts with Large Language Models (Lai et al., 2023; Yang et al., 2023; Rasheed et al., 2023), as shown in Fig. 1. The experimental results demonstrate that (1) using a multimodal encoder with the input text and image and (2) early fusion between the text and image contribute to the better-referring ability for SAM, showing a promising direction for text-prompted SAM. Additionally, the experiments also show the superiority of our EVF-SAM using a multimodal encoder over previous methods with decoder-only Large Language Models: (1) EVF-SAM reduces huge amounts of parameters, e.g., 82% parameters compared to LISA; (2) EVF-SAM relies less on handcrafted templates or instructions, which is more efficient and flexible; (3) EVF-SAM obtains better performance with less training data.
@@
  • We investigate the most effective approach to prompt SAM with texts by leveraging the Multimodal Encoder with multimodal inputs and the early vision-language fusion, which outperforms vanilla text encoders or Large Language Models.
  • We formulate the paradigm for text-prompted SAM and propose EVF-SAM, which is modular and readily integrated with mainstream foundation models. In addition, EVF-SAM gets rid of hand-crafted templates, and the training is stable and efficient compared to methods using Large Language Models.
  • The proposed EVF-SAM, only trained with open-source datasets, achieves state-of-the-art performance on the referring expression segmentation tasks, i.e., RefCOCO/+/g, demonstrating the effectiveness of our paradigm. Notably, EVF-SAM reduces parameters by 82% (1.3B v.s. 7.7B) compared to previous works based on Large Language Models.
@@ 2.1 Text-Prompted Segment Anything Models
Segment Anything Model. SAM (Kirillov et al., 2023) is an interactive segmentation model capable of predicting non-semantic masks based on various types of prompts (points, boxes, coarse masks). Trained on a large-scale dataset, SAM demonstrates strong generalization capability for segmenting diverse common objects. Several works (Xiong et al., 2023; Zhao et al., 2023) address the massive computation cost of SAM and propose efficient variants. Efficient-SAM (Xiong et al., 2023) distils the image encoder of SAM, achieving comparable performance with significantly fewer parameters. Fast-SAM (Zhao et al., 2023), leveraging the YOLOv8 (Jocher et al., 2023) architecture, achieves a 50×50$\times$50 × speedup for inference. SAM-HQ (Ke et al., 2024) addresses the segmentation quality of SAM and utilizes low-level features from the image encoder to enhance the mask decoder for better accuracy. Although SAM excels in visual-based segmentation tasks with box/point/mask prompts, it currently lacks language understanding abilities and it’s infeasible to directly use text prompts for referring segmentation or semantic segmentation.
Text-Prompted explorations. Recently, several works (Ren et al., 2024; Zhao et al., 2023; Li et al., 2023) have explored text prompts for SAM to segment objects according to the instructions or referring expressions. Grounded-SAM (Ren et al., 2024) leverages the Grounding DINO (Liu et al., 2023d) to obtain text-prompted boxes and feed the boxes to SAM for segmentation results, which formulates the non-end-to-end two-stage frameworks. Fast-SAM (Zhao et al., 2023) matches the similarity of CLIP (Radford et al., 2021) features between the text and Region of Interest (RoI) of image. RefSAM (Li et al., 2023) employs a lightweight cross-modal MLP to project the text embeddings of the referring expressions into SAM’s sparse embeddings and dense embeddings. LISA (Lai et al., 2023; Yang et al., 2023) employs a Large Multimodal Model, e.g., LLaVA (Liu et al., 2023b) to extract multimodal embeddings for SAM through the auto-regressive decoder. The aforementioned methods either suffer from poor performance or are computationally expensive. Referring expression segmentation based on SAM is a promising area for exploration, offering significant potential. We propose an effective end-to-end model that overcomes SAM’s limitations by enabling text-prompted segmentation capabilities.
@@ 2.2 Referring Expression Segmentation
Referring Segmentation via Text Encoders. Prevalent methods (Li & Sigal, 2021b; Wang et al., 2022b; Yang et al., 2022; Liu et al., 2023c) tend to leverage transformer-based text encoders, e.g., BERT (Devlin et al., 2018) or CLIP (Radford et al., 2021), to encode expression texts into embeddings as guidance for segmentation. RefTr (Li & Sigal, 2021a) uses a visual-language encoder to fuse image and text features and regresses the box and mask with a carefully designed query processor. LAVT (Yang et al., 2022) leverages a hierarchical Vision Transformer (Dosovitskiy et al., 2020) (ViT) to perform language-aware visual encoding. CRIS (Wang et al., 2022b) designs a vision-language decoder to merge CLIP features, propagating fine-grained semantic information from textual representations to each pixel-level activation. PolyFormer (Liu et al., 2023c) follows the encoder-decoder structure, employing a transformer decoder to generate regression results. Novel methods pay attention to being compatible with multiple tasks to formulate a uniform model. UNINEXT (Yan et al., 2023), UniRef++ (Wu et al., 2023) and UniLSeg (Liu et al., 2023e) employ similar frameworks but focus on utilizing datasets from different fields to empower their generalization capability. Although these traditional models are usually lightweight and achieve fine performance, They fail to integrate with large-scale foundation models, e.g., SAM(Kirillov et al., 2023), LLaVA(Liu et al., 2023b), thereby struggling to keep pace with the trend of increasingly extensive pre-training.
@@
Referring Segmentation via Large Language Models. In the context of the rapid development of Large Multimodal Models (Liu et al., 2023b; Bai et al., 2023; Sun et al., 2023b; a) (LMM), a number of works (Lai et al., 2023; Yang et al., 2023; Ren et al., 2023; Pi et al., 2023; Xu et al., 2023; Zhang et al., 2024) have leveraged these models to encode expression texts for referring expression segmentation tasks. LISA (Lai et al., 2023; Yang et al., 2023) finetune LLaVA (Liu et al., 2023b) to make it able to answer questions related to segmentation with a fixed template like ‘It is [SEG].’, where the hidden embeddings at the place of special token [SEG] will be seen as multimodal features extracted by LMM. PixelLM (Ren et al., 2023) extends LISA by building a segmentation codebook to enable multi-object segmentation. PixelLLM (Xu et al., 2024) empowers the vision-language model to take locations (e.g., a set of points or boxes) as either inputs or outputs. PerceptionGPT (Pi et al., 2023) proposes an end-to-end architecture. u-LLaVA (Xu et al., 2023) supports multi-task. PSALM (Zhang et al., 2024) imports mask tokens to LMM input for better performance. However, those methods tend to adopt heavy architectures, especially the LLMs or LMMs, leading to a heavy computation burden for downstream applications. In contrast, we find that lightweight vision-language models perform better for encode text prompts for referring image segmentation.
3 Method
3.1 Motivation: SAM with Vision-Language Models
Figure 3: Architectural explorations for text-prompted SAM. ‘L’ and ‘V’ denote the text encoder and vision encoder. We mainly explore three schemes: (a) vanilla baseline with a simple text encoder, (b) multimodal inputs with a late fusion, i.e., concatenation, and (c) multimodal inputs with early vision-language fusions, e.g., BEIT-3 (Wang et al., 2022a).
Considering that SAM (Kirillov et al., 2023) has a strong generalization capability for image segmentation while the text-prompted ability has not been revealed, we investigate how to encode text prompts for SAM in this section. We started by using the vanilla text encoder, as shown in Fig. 3 and conducted preliminary experiments on RefCOCO (testA) to evaluate the referring ability of SAM, shown in Tab. 1.
Multimodal referring information for SAM. SAM (Kirillov et al., 2023) has explored the feasibility of employing a CLIP text encoder to facilitate text-prompted segmentation, as illustrated in Fig. 3 (a). We owe its weak performance to the single-modal referring information. CLIP-prompted SAM achieves 63.4 cIoU at the RefCOCO/testA benchmark, far from well-defined baselines. CLIP exhibits strong alignment between text and image modalities, this alignment is insufficient for fine-grained tasks like segmentation. The referring information extractor should be provided with the input image and text prompts to ensure accurate alignment between the text expression and the relevant image region. We observe performance improvements after using multimodal prompts, i.e., 63.4 v.s. 67.9 for CLIP and 65.1 v.s. 83.7 for BEIT-3.
Table 1: Motivation analysis. Both CLIP and BEIT-3 are of Large scale, with comparable numbers of parameters. Specifically, CLIP has a total parameter count of 428M, while BEIT-3 totals 673M parameters. Metric of LLaVA (Liu et al., 2023b) is borrowed from LISA-7B (Lai et al., 2023)
CLIP (Text) CLIP (Text+Image) BEIT-3 (Text) BEIT-3 (Text+Image) LLaVA (Text+Image)
cIoU (RefCOCO) 63.4 67.9 65.1 83.7 79.1
Early-fused architecture. Some existing works, e.g., LISA (Lai et al., 2023; Yang et al., 2023), UniLSeg (Liu et al., 2023e), advocate for fusing visual and textual information simply before the mask generator and are widely considered as ‘early fusion’. However, we argue that these approaches are not early enough. As illustrated in Fig. 3 (b), we define such fusion for separately encoded single-modal prompts as ‘late fusion’. In contrast, as shown in Fig. 3 (c), we define the fusion during feature extraction, where both modalities can access the dense information of the other one, as ‘early fusion’, e.g., ViLT (Kim et al., 2021), BEIT-3 (Wang et al., 2022a), which incorporate the cross-modal fusions within the encoder.
We leverage the ‘early-fusion’ vision-language model as the Multimodal Encoder to generate prompt embeddings for SAM. Tab. 1 shows that our investigation indicates that early-fusion outperforms late-fusion, i.e., 83.7 for BEIT-3 and 67.9 for CLIP. We believe the early-fused architecture, as defined by our approach, is beneficial for encoding text prompts since the cross-modal fusions will further enhance the semantic representation for text embeddings. In addition, the text-to-image fusions guide the image branch to aggregate features which are aligned with text prompts, making the output embeddings more accurate for prompt SAM.
Encoder-based feature extractor. Recently, LISA (Lai et al., 2023; Yang et al., 2023) and several LLM-based methods (Rasheed et al., 2023; Ren et al., 2023) acquire the prompt embeddings for SAM with a special token through the auto-regressive generation. However, the uncontrolled length of the answering query introduces instability during both training and inference. Forcing the model to conform to a specific answering template can lead to language drift. In contrast, encoder-based architectures can maintain a consistent sequence length of inputs and outputs. Utilizing the encoder-based method not only offers convenience but also yields superior performance, i.e., 79.1 for LLaVA and 83.7 for BEIT-3. Notably, the encoder-based text-prompted SAM will reduce a massive computation burden compared to the LLM-based methods.
3.2 Architecture
Multimodal Encoder. The Early Vision-Language Fused encoder adopts the input image and text and outputs fused multimodal embeddings. In EVF-SAM, we mainly adopt BEIT-3 (Wang et al., 2022a) as the Multimodal Encoder, which formulates a multi-way transformer. The text is tokenized by XLMRobertaTokenizer (Conneau et al., 2019) while the image is resized to $224^{2}$ and patched by a 1/16 convolution layer. Within each block of the encoder, the image and text tokens will be fused in the attention block and then fed into separate Feed-Forward Networks (FFN). We follow ViT (Dosovitskiy et al., 2020) to retrieve the [CLS] token as the output multimodal embeddings.
Projector. Different foundation models tend to have different embedding dimensions (1024 for BEIT-3-Large, 768 for BEIT-3-Base, and 256 for SAM mask decoder). We adopt a simple MLP projector containing 2 Linear layers, activated by ReLU. In EVF-SAM, we do not design elaborate modules for better performance due to the following reasons: (1) the simple MLP is effective enough (Liu et al., 2023b; Kim et al., 2021), (2) using MLP is efficient for training and inference, and (3) the simple projector will have few impacts on the pre-trained knowledge of foundation models.
Adapted prompt encoder for SAM. SAM contains 3 main modules: (a) Image Encoder: a Vision Transformer (Li et al., 2022) (ViT), extracting fine-grained feature maps from the input image. (b) Prompt Encoder: receiving interactive prompts and encoding them into hidden embeddings. (c) Mask Decoder: a lightweight mask generator to output the final masks based on previous embeddings. In EVF-SAM, we maintain the architecture of the image encoder and mask decoder while extending the prompt encoder to further gather the embeddings from the Multimodal Encoder. Specifically, the original prompt encoder encodes point or box prompts to sparse embeddings of $R^{B\times N\times D}$, where $B$, $N$, and $D$ refer to the batch size, number of points/boxes, and the embedding dimension, respectively. In EVF-SAM, the projected multimodal embeddings of $R^{B\times 1\times D}$ from the Multimodal Encoder will be concatenated to a zero-initialized sparse embeddings and then fed into the mask decoder.
3.3 Training
Instruction template-free. In most LLM-based frameworks, e.g., LISA (Lai et al., 2023; Yang et al., 2023), instruction templates are required to prompt the Large Multimodal Models (LMM) for the segmentation task, e.g., ‘Can you segment {object} in the picture’ with answer ‘It is [SEG].’. Removing instruction templates will affect the performance of LMMs, therefore, users need to follow the corresponding templates for referring image segmentation. In contrast, EVF-SAM does not require pre-training on QA (question-answering) datasets, thus eliminating the need for instruction templates. We adopt the expression phrases or sentences as input. This template-free approach simplifies training and inference.
Trainable modules. The Multimodal Encoder (EVF) is fully trainable during our training process, allowing it to learn how to generate multimodal embeddings tailored for SAM, which requires sufficient localization information for segmentation. For SAM, we keep the image encoder frozen during training while we enable training for the prompt encoder and mask decoder. Our experiments revealed that freezing the prompt encoder and mask decoder only leads to a minimal performance drop while maintaining SAM’s ability. We present details in Sec. 3.
Unified training with multi-tasks. To further enhance the generic multi-task segmentation capabilities of EVF-SAM, including semantic segmentation and fine-grained part segmentation, we present a unified training strategy for EVF-SAM with diverse training datasets, such as ADE20K (Zhou et al., 2017b), PartImageNet (He et al., 2022) and PASCAL-Part (Chen et al., 2014).
However, we observe a performance degradation when simply mixing the training data of referring and semantic segmentation (shown in Tab. 7 of the Appendix), which can be attributed to the semantic conflict among different tasks, as discussed in UniLSeg (Liu et al., 2023e). To alleviate the aforementioned conflicts, we leverage a special text token [semantic] and input ‘[semantic]{category}’ for semantic/part segmentation.
4 Experiments
4.1 Datasets and Metrics
Datasets. We mainly conduct the experiments on RefCLEF (Kazemzadeh et al., 2014), RefCOCO, RefCOCO+ (Yu et al., 2016; Kazemzadeh et al., 2014), and RefCOCOg (Nagaraja et al., 2016; Mao et al., 2016). Specifically, RefCOCOg contains longer expressions which are manually annotated. Except for RefCOCO+, all datasets include geometric expression (e.g., ‘on the left’). Among different splits of testing datasets, ‘testA’ is human-centric, while ‘testB’ aims for common objects.
Extra training datasets. To further enhance the versatility of EVF-SAM, we employ multi-task unified training by expanding the training datasets by introducing Objects365 (Shao et al., 2019), ADE20K (Zhou et al., 2017b), PASCAL-Part (Chen et al., 2014), PartImageNet (He et al., 2022), and HumanParsing (Liang et al., 2015b). Therefore, EVF-SAM can handle various granularity of text-prompted segmentation, e.g., semantic-level, instance-level, and part-level segmentation. We refer the readers to the appendix for more details about training with the extra multi-task datasets.
4.2 Implementation Details
Unless specified, we initialize the proposed EVF-SAM with the public weights of SAM-ViT-Huge  [^1] SAM: https://github.com/facebookresearch/segment-anything (Kirillov et al., 2023) and BEIT-3-Large  [^2] BEIT-3: https://github.com/microsoft/unilm (Wang et al., 2022a). All models are trained on 4 NVIDIA L40s GPUs with mixed precision. We adopt DeepSpeed (Song et al., 2023) with ZeRO-2 for model parallel to optimize memory consumption. During training, the batch size of each GPU is 16 and we use gradient accumulation for 2 steps, therefore the total batch size per iteration is 128. We adopt AdamW (Loshchilov & Hutter, 2017) optimizer and set the initial learning rate to 1e-4 with a linear-decay schedule. We train all models for 15k iterations (nearly 1 day) and use the binary cross-entropy loss (BCE) and dice loss (the weight of both losses is 1.0).

### v5

@@ Abstract
Segment Anything Model (SAM) has attracted widespread attention for its superior interactive segmentation capabilities. Despite this, the potential of text prompts for SAM remains largely unexplored. In this paper, we investigate what text prompt encoders are beneficial for adapting SAM for Referring Expression Segmentation (RES) and observe that (1) multimodal prompts and (2) the early-fusion mechanism incorporating text-to-image attention are crucial factors. We introduce the Early Vision-Language Fusion (EVF) framework and further formulate the EVF-SAM, which is simple but demonstrates superior effectiveness and efficiency compared to mainstream methods based on Large Language Models (LLM). We also successfully integrate segmentation data from diverse tasks into an unified hybrid dataset to conduct joint training. The well-designed data strategies resolve issues such as semantic conflict and ambiguity. The proposed EVF-SAM based on BEIT-3 obtains state-of-the-art performance on RES tasks (79.0 in average on RefCOCO/+/g), and extends RES capability to various granularities, e.g., semantic-level RES and part-level RES. EVF-SAM maintains the model size with only 1.32B parameters, considerably fewer than the mainstream LLM-based methods of over 7B parameters. Code and models will be made publicly available.
@@ 1 Introduction
Segment Anything Model (SAM) [24] brings interactive segmentation paradigm to public view. Well-trained on the SA-1B dataset, SAM achieves stunning performance and quickly becomes popular as a vision foundation model for object localization and beyond. Various SAM variants [64, 74, 76, 22] have been explored, achieving better efficiency or higher precision. Despite SAM’s surprising abilities like point-prompted and box-prompted segmentation, it is a pity that the text-prompted segmentation ability remains conceptual. We retrospect this task to Referring Expression Segmentation (RES). RES focuses on the solution to predicting the segmentation mask according to the text description given by users, which enjoys several explorations by some traditional models [13, 34, 52, 4, 71, 14, 9, 27, 60, 69, 38, 61, 39, 67], and is broadened by some Large Language Models (LLM) [25, 68, 50, 44, 65, 75, 62, 47].
@@
Figure 2: Comparison between late-fusion and early-fusion. We visualize the attention map of representative late-fusion method (i.e., LISA [25]) and early-fusion methods (i.e., EVF-SAM) among different layers. The Y axis represents the ‘Query’, the X axis represents the ‘Key’. We use red, blue and orange boxes to highlight the ‘image to text attention’, ‘text-to-image attention’, and targeted segmentation object. The figures demonstrate that the ‘text-to-image attention’ is more crucial for Referring Expression Segmentation task, while late-fusion methods ignore it. To compensate for this, we propose the early-fusion method for high-quality text-guided visual features.
The key challenge lies in empowering SAM with language understanding ability for segmentation according to text prompts, e.g., referring expression segmentation. Fig. 3 summarizes previous end-to-end works which explore the text-prompted abilities of SAM: (a) SAM with text encoder: leverage an off-the-shelf text encoder(e.g., CLIP [45]) to extract text features and feed into SAM. However, text encoders fail to extract positional information from unimodal input, leading to poor performance. (b) SAM with LLM: fine-tune a Large Language Model (LLM) to auto-regressively generate embeddings for the target object. These methods are under the late-fusion approach, where the image features extracted by the image encoder are further integrated within the LLM. However, the text-to-image attention, which is proved essential (in Fig. 2) for Referring Expression Segmentation task, is neglected in LLMs. Additionally, these LLM-based models are often computationally expensive and hard to train.
To overcome shortcomings of existing technical approaches, we propose: (1) multimodal prompts that incorporate both text and image outperform text-only prompts; (2) the early-fusion architectures, which incorporate text-to-image attention, exhibit significant advantages over text-only encoders and Large Language Models. Motivated by the above observations, we realize the text-promptd SAM using an Early Vision-Language Fusion encoder (EVF), and propose EVF-SAM, as shown in Fig. 3 (c). The proposed EVF-SAM is a simple yet effective framework built upon the off-the-shelf foundation models, comprising an well-pretrained Vision-Language Model (VLM) (i.e., BEIT-3 [59]), to generate prompt embeddings for SAM. Besides, EVF-SAM does not include elaborate designs or modules and is easy for scaling to larger models. Moreover, EVF-SAM reduces huge amounts of parameters (e.g., -82% parameters compared to LISA [25]), with better performance outperforming previous attempts with Large Language Models [25, 68, 47], as shown in Fig. 1.
Besides the simple architecture, we highlight the importance of well-managed training data. We construct a hybrid dataset with publicly available datasets by including RES data (e.g., RefCOCO/+/g [21, 72, 42, 41]), semantic-level data (e.g., ADE20K [77, 78]), instance-level data (e.g., Objects365 [51]), and part-level data (e.g., Pascal-Part [5]). To effectively train the model on the diverse collection of hybrid dataset, we carefully design several data strategies for our EVF-SAM. To be specific, we propose to use a special token ‘[semantic]’ to solve the semantic conflict problem, and use some merging/filtering principles to solve the ambiguity problem. These additional data not only boost EVF-SAM at Referring Expression Segmentation (RES) tasks (78.0 cIoU to 79.0 cIoU in average on RefCOCO/+/g [21, 72, 42, 41]), but also extend EVF-SAM’s capabilities at various granularity (e.g., semantic-level RES, part-level RES).
@@
  • We propose a novel Referring Expression Segmentation (RES) framework by introducing an Early Vision-Language Fusion encoder (EVF). The EVF framework demonstrates superior effectiveness and efficiency compared to mainstream late-fusion frameworks, e.g., RES via Large Language Models (LLM). The EVF framework emphasizes the multimodal prompt and the early-fusion mechanism, which are crucial for the RES task, as highlighted by extensive experiments and visualizations. Building upon the EVF framework, we propose EVF-SAM.
  • We jointly train EVF-SAM on a hybrid dataset, extending the model’s capability to various granularities of RES tasks, e.g., semantic-level RES, part-level RES. Our proposed data strategies successfully map various segmentation datasets to a unified representation, resolving distribution problems such as semantic conflict and ambiguity.
  • EVF-SAM achieves state-of-the-art performance on the RES tasks (i.e., 79.0 cIoU in average on RefCOCO/+/g). Meanwhile, EVF-SAM reduces the parameter count to 1.3B, a substantial decrease compared to the LLM-based models of more than 7B parameters.
Figure 3: Comparisons of different Text-prompted SAM. (a) A natural idea to support text prompts is to use an off-the-shelf text encoder to generate text embeddings for SAM [24, 30]. (b) Several works [25, 68, 47] adopt Large Language Models (LLM) to generate prompt embeddings for SAM in an auto-regressive manner. (c) Our proposed EVF-SAM exploits an effective early vision-language fusion encoder for text-prompted SAM with higher performance and fewer parameters.
Figure 4: Architectural explorations for text-prompted SAM. ‘L’ and ‘V’ denote the text encoder and vision encoder. We mainly explore three schemes: (a) vanilla baseline with a simple text encoder, (b) multimodal inputs with a late-fusion, i.e., concatenation, and (c) multimodal inputs with early-fusion
@@ 2.1 Text-Prompted Segment Anything Models
Segment Anything Model. SAM [24] is an interactive segmentation model capable of predicting masks based on various types of prompts (points, boxes, and masks). Trained on a large-scale dataset, SAM demonstrates strong generalization capability for segmenting diverse common objects. Advanced works address the massive computation cost of SAM or broaden SAM’s capability. Efficient-SAM [64] distills the image encoder of SAM, achieving comparable performance with significantly fewer parameters. Fast-SAM [76], leveraging the YOLOv8 [18] architecture, achieves a 50×50$\times$50 × speedup for inference. SAM-HQ [22] utilizes low-level features from the image encoder to address the segmentation quality of SAM. Matte-Anything [70] make SAM capable of outputting detailed matting masks. Inpaint-Anything [73] integrate SAM with AIGC models to produce objects-removing tools. VRP-SAM [56] enables visual prompt for referring segmentation. Edit Everything [63] allows users to edit images using simple text instructions.
Text-prompted SAM. Although SAM excels in visual-based segmentation tasks with box/point/mask prompts, it currently lacks language understanding abilities and it’s infeasible to directly use text prompts for Referring Expression Segmentation (RES). Several works [49, 76, 30] employ multi-stage architectures (e.g., employing grounded detectors) to incorporate SAM with language interpretation capability. However, multi-stage models suffer from noise accumulation and difficulty in optimization, and is out of our concern. Under end-to-end manners, vanilla SAM proposes employing a CLIP text encoder to cast linguistic feature into SAM, as shown in Fig. 4 (a). Some advanced works propose using late-fusion manner, by employing off-the-shelf unimodal encoders to pre-extract features independently, followed by a fusion structure, as shown in Fig. 4 (b). RefSAM [30] employs a lightweight cross-modal MLP to project the text embeddings into SAM’s prompt encoder LISA [25, 68] employs a LLM (LLaVA [36]) to extract multimodal embeddings for SAM through the auto-regressive decoder. The aforementioned methods either suffer from poor performance or are computationally expensive. In this work, our EVF-SAM is in ‘early-fusion’ manner, which is both effective and efficient, as shown in Fig. 4 (c).
@@ 2.2 Referring Expression Segmentation
RES via Text Encoders. Prevalent methods tend to leverage transformer-based text encoders (e.g., BERT [7], CLIP [45]) to encode texts into embeddings as guidance for segmentation. RefTr [26] uses a vision-language encoder to fuse multimodal features and regresses the box and mask with a carefully designed query processor. LAVT [69] leverages a hierarchical ViT [10] to perform language-aware visual encoding. CRIS [60] designs a vision-language decoder to merge CLIP features, propagating fine-grained semantic information from textual representations to each pixel-level activation. PolyFormer [38] follows the encoder-decoder structure, employing a transformer decoder to generate regression results. Novel methods pay attention to being compatible with multiple tasks to formulate a uniform model. UNINEXT [67], UniRef++ [61] and UniLSeg [39] employ similar frameworks but focus on utilizing datasets from different fields to empower their generalization capability. Although these traditional models are usually lightweight and achieve fine performance, they fail to integrate with large-scale foundation models(e.g., SAM[24], LLaVA[36]), thereby struggling to keep pace with the trend of increasingly extensive pre-training.
@@
RES via Large Language Models. In the context of the rapid development of Large Language Models [36, 1, 55, 54] (LLM), a number of works [25, 68, 50, 44, 65, 75] have leveraged these models to encode expression texts for referring expression segmentation tasks. LISA [25, 68] finetunes LLaVA [36] to make it answer questions related to segmentation with a fixed template like ‘It is [SEG].’, where the hidden embeddings at the place of special token [SEG] will be seen as multimodal features. PixelLM [50] extends LISA by building a segmentation codebook to enable multi-object segmentation. PixelLLM [66] empowers the vision-language model to take locations (e.g., a set of points or boxes) as either inputs or outputs. PerceptionGPT [44] proposes an end-to-end architecture by designing a perception-enhanced vision language model, which improves inference efficiency. u-LLaVA [65] supports multi-task, e.g., object grounding. PSALM [75] imports mask tokens to LLM input for better performance. However, LLMs neglect the text-to-image attention, which is proved essential for RES in our later experiments, leading to their sub-optimal performance. Additionally, they tend to adopt heavy architectures, especially the LLMs, leading to a heavy computation burden for downstream applications. In contrast, we find that lightweight early vision-language fusion models perform better for encode text prompts for referring image segmentation.
2.3 Late-fusion and Early-fusion
Late-fusion models. Late-fusion denotes applying off-the-shelf unimodal encoders(e.g., CLIP [45]) to pre-extract features independently, followed by a fusion structure. This leverages the strong prior knowledge gained during the encoders’ pre-training, requiring only minimal adjustment (e.g., simple projection layers) to align the multimodal feature space. This is particularly advantageous for models with complex mechanisms (e.g., causal attention), easing the convergence and gain the performance. Representative late-fusion methods are Multimodal Large Language Models (MLLM). LLaVA [36]) employs the off-the-shelf CLIP image encoder (CLIP [45]) to pre-extract visual features. The visual features is projected to a Large Language Model (i.e., LLaMA [57]) using a simple linear layer. Furthermore, Qwen-VL [1] utilizes finer-designed training recipe and carefully managed data, building a robust MLLM. Mipha [79] explores the lightweight architecture, following LLaVA’s recipe to finetune phi [29]. There are also attempts attempts to decouple causal models from late fusion, e.g., Fuyu [2], EVE [8]. But they exhibit significantly slower convergence and poorer performance compared to late-fusion MLLMs.
Early-fusion models. Early-fusion methods proposes using shallow unimodal embedding layers to project features, and then using deep fusion layers to conduct multimodal interaction. The early integration of fusion make both modalities attach to dense information of the other one, so the extraction of multimodal feature is advantageously effective and efficient. This is crucial for tasks strongly relied on the enhance of all included modalities. Representative early-fusion methods are Vision-Language Models (VLM) with bidirectional attention. ViLT [23] conducts detailed comparisons to prove the effectiveness and efficiency of early-fusion methods over late-fusion methods (e.g., Pixel-BERT [15]) in terms of multimodal feature extraction. BEIT-3 [59] project visual signal to language space and performs masked language modeling pretraining, building a powerful backbone. ONE-PEACE[58] extends this approach to incorporate audio modalities, along with utilizing a larger volume of pre-training data and increasing the model’s parameter size.
In this work, we proposed that ‘early-fusion’ (i.e., EVF-SAM) is prioritized over ‘late-fusion’ (i.e., ‘RES via text encoder’, ‘RES via LLMs’, introduced in Sec. 2.2) in the Referring Expression Segmentation task.